OpenStudy (anonymous):

Find an interval of length 0.1 in [1, 2] containing a root of the equation x^7+4 x-10=0.

OpenStudy (amistre64):

i wonder what we get if we divide it by x-1 and x-2

OpenStudy (amistre64):

otherwise its just an interative process of your choosing

OpenStudy (amistre64):

dividing doesnt seem to have any practical applications ... it was just a thought

OpenStudy (amistre64):

newton had a method of taking the derivative and recaluclating

myininaya (myininaya):

lol sorry i was thinking integrals for a second

OpenStudy (amistre64):

the linear approximation of f(x) at c is: y - f(c) = f'(c) (x-c) finding the x intercept gives us a new guess to play with \ xn = -f(c)/f'(c) + cf'(c)

OpenStudy (amistre64):

when x_(n+1) - x_n <= .1/2 then we have found one of the endpoints of the interval maybe doesnt need to be /2 ... but yeah